Grace bought an 800-peso jacket at 20% discount and a 1,200-peso shoes at a 10% discount. how much did she save?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39615)
Answer by ikleyn(52770)

.

She saved  0.2*800 + 0.1*1200 peso = 160 + 120 = 280 peso.